# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

### v2

@@ 4 Stability-driven teacher model selection
In Section 3, we showed that the benefits of distillation arise from the teacher model’s ability to construct a meaningful basis representation of the individual-level treatment effect using the covariate data. In practice, there are different metalearners that researchers must choose from to estimate the individual-level treatment effect. Traditional model selection procedures cannot be used to evaluate the goodness-of-fit of our teacher models, which predict individual-level treatment effects — quantities that are not observable in practice. Furthermore, checking whether the predicted outcomes match the observed outcomes is insufficient to guarantee that the treatment effects are modeled well (Künzel et al., 2019).
We next introduce a novel model selection procedure, designed specifically for subgroup estimation. Our proposed procedure selects a teacher model based on its ability to stably reconstruct subgroups. Stable reconstruction refers to being able to estimate similar subgroups given various perturbations in the underlying data generating process. We consider more stable teacher models as being more appropriate models for the data, as they demonstrate greater robustness to sampling error (e.g., Yu, 2013; Yu & Kumbier, 2020).
@@ 5 Simulations
Figure 3: Performance of subgroup estimation methods for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via RMSE for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes with linear covariate effects (rows). Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
@@ 6 Case Study: AIDS Clinical Trials Group Study 175
These three covariates interact across our subgroups, producing varying treatment effects. Those with a higher starting weight benefit more from combination therapy under both lower and higher CD8 counts (i.e., Groups 2 and 3 vs Group 1), though the effect is slightly stronger under a lower (healthier) CD8 count. Notably, it is only when those with a lower starting weight also have the highly risky combination of high CD8 cell counts and substantial previous therapy 666786 days of therapy is in the upper 78th percentile of previous exposure in this study that they benefit significantly from the treatment (Group 4).
@@ 7 Discussion
We introduced causal distillation trees for interpretable subgroup estimation in causal inference. CDT allows researchers to leverage the flexibility and power of black-box metalearner approaches, while preserving the interpretability of simple decision trees. We prove that CDT allows researchers to consistently estimate the optimal set of subgroups, and we introduce a novel model selection procedure to help researchers use CDT in practice. Furthermore, we have developed an R package causalDT available at https://github.com/tiffanymtang/causalDT, which not only fits CDT but also automatically conducts the subgroup stability and diagnostic checks, thus enabling researchers to interpret their estimated subgroups from CDT alongside important and informative diagnostics in a holistic manner.
There are several promising avenues for future research. First, recent literature in external validity has emphasized the importance of recruiting a sufficiently heterogenous experimental sample that adequately captures the total treatment effect variation across the target population (e.g., Huang & Parikh, 2024; Huang, 2024+). CDT offers an opportunity for researchers to characterize the underlying treatment effect heterogeneity in a given study and to consider how to optimally recruit these subgroups of participants for an externally valid study. Second, CDT provides a very general and flexible two-stage framework. While we focused on CART as the second-stage learner, there are numerous alternative tree-based models (e.g., Loh, 2002; Quinlan, 2014; Hu et al., 2019) and many that have been developed specifically for subgroup estimation (e.g., Loh et al., 2015; Su et al., 2009; Lipkovich et al., 2011; Seibold et al., 2016; Wang et al., 2024), which can be used as second-stage learners in CDT and may further improve performance. Similarly, efforts to further enhance the stability of distillation-based methods (Zhou et al., 2024) can be easily incorporated into the broader CDT framework. Lastly, while this work is focused on the causal inference setting and treatment effect heterogeneity, an extension of CDT could consider prediction models, where there may be variation in the underlying predictive performance of a specific model. In this prediction context, subgroup analyses could help diagnose critical disparities in the performance of predictive models.
@@ A.2 Extended discussion on valid teacher models
Let $s$ and $s^{d}$ respectively be the optimal splits using the original individual-level treatment effects $\tau_{i}$ and predicted individual-level treatment effects $\hat{\tau}_{i}^{d}$, as defined in Assumption (3.2). Also, define $v_{i}:=\tau_{i}-\hat{\tau}_{i}^{d}$ for each $i=1,\ldots,n$. We can then write
@@ A.3 Inference for observational settings
A.4 Extension to infinite population model
In the main manuscript, we focus on the finite-sample setting for considering inference for the subgroup difference-in-means estimator. In settings when researchers are interested in estimating the population-level subgroup ATE, if the sample is a random draw from this larger target population, then we can show that the subgroup difference-in-means estimator will be a consistent estimator. To begin, let $\mathcal{D}$ represent the finite-sample study data. Then:
Let $K_{\tau}=\max_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left\{Y_{i}(1)-Y_{i}(0)\right\}$:
$\leq\tau^{(g)}+\frac{1}{n_{g}}\sum_{i=1}^{n}K_{\tau}\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert$
The variance of the subgroup difference-in-means estimator would look similar to the variance expression in Theorem 3.3, but with the population-level parameters. Furthermore, when considering the population-level inference, the cross correlation of potential outcomes will disappear. As a result, while the variance estimator will be the same for both inference in a finite-sample setting and an infinite super-population setting, we generally expect more uncertainty associated with inference for the population-level subgroup ATE.
If the study is not a random sample from target population, then we must account for potential distribution shifts between the study sample and the target population. Being able to transport subgroups across different studies to different target populations would require stronger assumptions (see, for example, Zhang et al. 2024).
@@ A.5 Non-parametric test of treatment effect heterogeneity
$\boldsymbol{\hat{\tau}}=(\hat{\tau}^{(1)}-\hat{\tau},\ldots,\hat{\tau}^{(g)}-\hat{\tau})^{\top},$
where $\hat{\tau}:=\frac{1}{\sum_{i=1}^{n}Z_{i}}Y_{i}Z_{i}-\frac{1}{\sum_{i=1}^{n}(1-Z_{i})}Y_{i}(1-Z_{i})$ is the usual difference-in-means estimator across the entire sample. Then, following Theorem 3.3, we can show that under the null hypothesis, the test statistic can be asymptotically approximated using a $\chi^{2}$ distribution. The following corollary formalizes.
@@
$\boldsymbol{\hat{\tau}}^{\top}\boldsymbol{\Sigma}\boldsymbol{\hat{\tau}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\chi^{2}.$
@@ A.6 Detailed CDT Algorithm
2 // Stage 1: estimate individual-level treatment effects via teacher model
@@ A.7 Discussion of Teacher Model Selection Algorithm
2      Fit $\mathcal{M}^{l}$ on $\mathcal{D}^{train}$ to obtain estimates of the individual-level treatment effects $\hat{\tau}^{d,l}_{i}$ for each unit in $\mathcal{D}^{train}$ (via lines 2-16 in Algorithm 1)
@@ Appendix B Proofs and Derivations
×$\qquad\times\Pr\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}_{n}^{(a)})\right\}<\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}_{n}^{(b)})\right\}\right]$
@@
B.3 Proof of Theorem 3.2
To prove Theorem 3.2, we will introduce a lemma that will be helpful in the following derivation. We define $n_{g}$ as the total number of units in a subgroup $g$, and $n_{g}(1)=\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})Z_{i}$ – i.e., $n_{g}(1)$ is equal to the number of units in subgroup $g$ who are treated. We similarly define $n_{g}(0)$ as the number of units in subgroup $g$ who are not treated.
@@
    Equiprobable treatment assignment patterns. For each subgroup $g\in\{1,\ldots,G\}$, all ($n_{g}\choose n_{g}(1)$ combination of ways to treat $n_{g}(1)$ units are equiprobable, given the subgroup size $n_{g}$.
@@
We are considering desings that satisfy assignment symmetry. For example, experimental designs such as complete randomization, Bernoulli randomization, or blocking with independent treatment assignment across blocks, would satisfy assignment symmetry. However, cluster randomization would not satisfy assignment symmetry. Furthermore, we do not account for settings in which some units within a subgroup are more likely to receive treatment.
@@
$\mathbb{E}\left\{Z_{i}\mid n_{g}(1)\right\}=\frac{n_{g}(1)}{n_{g}}\equiv\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})Z_{i},\text{ and }\mathbb{E}\left\{Z_{i}Z_{j}\mid n_{g}(1)=a\right\}=\frac{a(a-1)}{n_{g}(n_{g}-1)}$ (9)
@@
  • $\mathbb{E}\left[\frac{Z_{i}}{n_{g}(1)}\right]=\frac{1}{n_{g}}$
  • $\mathbb{E}\left[\frac{Z_{i}^{2}}{n_{g}(1)^{2}}\right]=\frac{1}{n_{g}}\mathbb{E}\left[\frac{1}{n_{g}(1)}\right]$
  • $\mathbb{E}\left[\frac{Z_{i}Z_{j}}{n_{g}(1)^{2}}\right]=\frac{1}{n_{g}(n_{g}-1)}\mathbb{E}\left[\frac{n_{g}(1)-1}{n_{g}(1)}\right]=\frac{1}{n_{g}(n_{g}-1)}\left(1-\mathbb{E}\left[\frac{1}{n_{g}(1)}\right]\right)$
  • $\mathbb{E}\left[\frac{Z_{i}(1-Z_{j})}{n_{g}(1)n_{g}(0)}\right]=\frac{1}{n_{g}\cdot(n_{g}-1)}\cdot\mathbb{E}\left[\frac{n_{g}-n_{g}(1)}{n_{g}(0)}\right]=\frac{1}{n_{g}\cdot(n_{g}-1)}$
@@
Part 2. Unbiasedness. To begin, we will show that $\hat{\tau}^{(g)}$ is unbiased for the within-sample subgroup ATE:
$\mathbb{E}$ $\left(\hat{\tau}^{(g)}\right)=\mathbb{E}\left\{\left.\frac{1}{\sum_{i=1}^{n}Z_{i}\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}Z_{i}Y_{i}\hat{\mathcal{G}}_{g}(X_{i})-\frac{1}{\sum_{i=1}^{n}(1-Z_{i})\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}(1-Z_{i})Y_{i}\hat{\mathcal{G}}_{g}(X_{i})~{}\right|\mathcal{D}\right\}$
@@
var $\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1)\right\}$ (12)
$=\sum_{i,j:\hat{\mathcal{G}}_{g}(X)=1}\text{cov}\left(\frac{Z_{i}}{n_{g}(1)}Y_{i}(1),\frac{Z_{j}}{n_{g}(1)}Y_{j}(1)\right)$
$=\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[\text{var}\left\{\frac{Z_{i}}{n_{g}(1)}Y_{i}(1)\right\}+\sum_{i\neq j}\text{cov}\left\{\frac{Z_{i}}{n_{g}(1)}Y_{i}(1),\frac{Z_{j}}{n_{g}(1)}Y_{j}(1)\right\}\right]$
$=\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[\frac{1}{n_{g}}\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}Y_{i}(1)^{2}-\frac{1}{n_{g}^{2}}Y_{i}(1)^{2}\right]+$
$\sum_{i\neq j}\frac{1}{n_{g}}\cdot\frac{1}{n_{g}-1}\cdot\left(1-\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}\right)Y_{i}(1)Y_{j}(1)-\frac{1}{n_{g}^{2}}Y_{i}(1)Y_{j}(1)$
$=\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left(\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}-\frac{1}{n_{g}}\right)Y_{i}(1)^{2}+$
$~{}~{}~{}~{}\frac{1}{n_{g}}\sum_{i\neq j:\hat{\mathcal{G}}_{g}(X)=1}\left(\frac{1}{n_{g}-1}\left(1-\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}\right)-\frac{1}{n_{g}}\right)Y_{i}(1)Y_{j}(1)$
Let $\tilde{\beta}_{g}(z):=\mathbb{E}\{1/n_{g}(1)\}$:
@@
=$=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\cdot\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[Y_{i}(1)^{2}-\sum_{i\neq j}Y_{i}(1)Y_{j}(1)\right]$
=$=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(1)\right\}$ (13)
We can similarly show term (2) can be rewritten as:
$\text{var}\left\{\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0)\right\}=\left\{\beta_{g}(0)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(0)\right\}.$ (14)
@@
$\mathbb{E}$ [$\left[\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1)\right\}\left\{\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0)\right\}\right]$
$=\mathbb{E}\left\{\frac{1}{n_{g}(1)\cdot n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}(1-Z_{i})Y_{i}(1)Y_{i}(0)+\sum_{i\neq j}Z_{i}(1-Z_{j})Y_{i}(1)Y_{j}(0)\right\}$
@@
$=\mathbb{E}\left\{\frac{1}{n_{g}(1)\cdot n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\sum_{i\neq j}Z_{i}(1-Z_{j})Y_{i}(1)Y_{j}(0)\right\}$
@@
cov $\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1),\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0),\right\}$

### v3

@@ 4 Stability-driven teacher model selection
In Section 3, we showed that the benefits of distillation arise from the teacher model’s ability to construct a meaningful basis representation of the individual-level treatment effect using the covariate data. While numerous metalearners have been developed and can be used as the teacher model, researchers may wonder how to choose an appropriate teacher model for CDT. Traditional model selection procedures cannot be used to evaluate the goodness-of-fit of our teacher models, which predict heterogeneous treatment effects — quantities that are not observable in practice. Furthermore, checking whether the predicted outcomes match the observed outcomes is insufficient to guarantee that the treatment effects are modeled well (Künzel et al., 2019).
Given these challenges, we propose a novel stability-based model selection procedure, designed specifically for subgroup estimation. Our proposed procedure selects a teacher model based on its ability to stably reconstruct subgroups. Stable reconstruction refers to being able to estimate similar subgroups given various perturbations in the underlying data generating process. We consider more stable teacher models as being more appropriate models for the data, as they demonstrate greater robustness to sampling error (e.g., Yu, 2013; Yu & Kumbier, 2020). Moreover, as we will see through simulations in Section 5, teacher models with higher stability generally correspond to more accurate estimation of the heterogeneous treatment effects and thus provides potential plausibility for the validity of the teacher model.
@@ 5 Simulations
Figure 3: Performance of subgroup estimation methods for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via RMSE for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes with linear covariate effects (rows). Results are averaged across 100 simulation replicates with ribbons denoting $\pm 1SE$.
@@ 6 Case Study: AIDS Clinical Trials Group Study 175
These three covariates interact across our subgroups, producing varying treatment effects. Those with a higher starting weight benefit more from combination therapy under both lower and higher CD8 counts (i.e., Groups 2 and 3 vs Group 1), though the effect is slightly stronger under a lower (healthier) CD8 count. Notably, it is only when those with a lower starting weight also have the highly risky combination of high CD8 cell counts and substantial previous therapy 777786 days of therapy is in the upper 78th percentile of previous exposure in this study that they benefit significantly from the treatment (Group 4).
@@ 7 Discussion
We introduced causal distillation trees for interpretable subgroup estimation in causal inference. CDT allows researchers to leverage the flexibility and power of black-box metalearner approaches, while preserving the interpretability of simple decision trees. We prove that CDT allows researchers to consistently estimate the optimal set of subgroups, and we introduce a novel model selection procedure to help researchers use CDT in practice. Furthermore, we have developed an R package causalDT which not only fits CDT but also automatically conducts the subgroup stability and diagnostic checks, thus enabling researchers to interpret their estimated subgroups from CDT alongside important and informative diagnostics in a holistic manner.
There are several promising avenues for future research. First, recent literature in external validity has emphasized the importance of recruiting a sufficiently heterogenous experimental sample that adequately captures the total treatment effect variation across the target population (e.g., Huang & Parikh, 2024; Huang, 2024+; Zhang et al., 2024). CDT offers an opportunity for researchers to characterize the underlying treatment effect heterogeneity in a given study and to consider how to optimally recruit these subgroups of participants for an externally valid study. Second, CDT provides a very general and flexible two-stage framework. While we focused on CART as the second-stage learner, there are numerous alternative tree-based models (e.g., Loh, 2002; Quinlan, 2014; Hu et al., 2019) and many that have been developed specifically for subgroup estimation (e.g., Loh et al., 2015; Su et al., 2009; Lipkovich et al., 2011; Seibold et al., 2016; Wang et al., 2024), which can be used as second-stage learners in CDT and may further improve performance. Similarly, efforts to further enhance the stability of distillation-based methods (Zhou et al., 2024) can be easily incorporated into the broader CDT framework. Lastly, while this work is focused on the causal inference setting and treatment effect heterogeneity, an extension of CDT could consider prediction models, where there may be variation in the underlying predictive performance of a specific model. In this prediction context, subgroup analyses could help diagnose critical disparities in the performance of predictive models.
@@ A.2 Extended discussion on valid teacher models
Let $s$ and $s^{d}$ respectively be the optimal splits using the original individual-level treatment effects $\tau_{i}$ and predicted heterogeneous treatment effects $\hat{\tau}_{i}^{d}$, as defined in Assumption (3.2). Also, define $v_{i}:=\tau_{i}-\hat{\tau}_{i}^{d}$ for each $i=1,\ldots,n$. We can then write
@@ A.3 Inference for observational settings
A.4 Extension to finite population model
In the main manuscript, we focus on the infinite target population setting for considering inference for the subgroup difference-in-means estimator. In settings when researchers are interested in estimating the finite population-level subgroup ATE, then we can show that the subgroup difference-in-means estimator will be unbiased for the sample subgroup ATE.
To begin, we define the finite-sample subgroup ATE as:
$\tau^{(g)}(\mathcal{D})=\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X)\{Y_{i}(1)-Y_{i}(0)\}.$
Let $\mathcal{D}$ represent the finite-sample study data.
Theorem A.1 (Unbiasedness and Variance of Subgroup Difference-in-Means Estimator in Finite Populations).
Under Assumptions (3.1)-(3.7), the subgroup difference-in-means estimator will be an unbiased estimator for the sample subgroup ATE:
$\mathbb{E}(\hat{\tau}(g)\mid\mathcal{D})=\tau^{(g)}(\mathcal{D}).$
The variance of $\hat{\tau}^{(g)}$ is
$\text{var}(\hat{\tau}^{(g)}\mid\mathcal{D})=\frac{1}{n_{g}}\left\{\beta_{g}(1)\text{var}_{g}(Y_{i}(1))+\beta_{g}(0)\text{var}_{g}(Y_{i}(0))-\text{var}_{g}(\tau_{i})\right\},$
[…]
To consider the asymptotic distribution of the subgroup difference-in-means estimator, we must account for the fact that as the sample size increases, the underlying subgroup allocation will also change. We leverage recent work in developing central limit theorem results for finite populations (e.g., Schochet 2024, Li & Ding 2017) to show that under mild regularity conditions, the subgroup difference-in-means estimator will be asymptotically normal.
Theorem A.2 (Asymptotic Normality of Subgroup Difference-in-Means Estimator).
Assume that as the sample size $n\to\infty$, the total number of subgroups $G$ remain fixed. Furthermore, assume the proportion of units in each subgroup converges to a proportion $\pi_{g}^{*}\in(0,1)$ (i.e., $n_{g}/n\to\pi_{g}^{*}$, where $0<\pi_{g}^{*}<1$ for all $g\in\{1,\ldots,G\}$ and $\sum_{g=1}^{G}\pi_{g}^{*}=1$), and the proportion of treated units converges to $p_{z}$, where $p_{z}\in(0,1)$ (i.e., $\frac{1}{n}\sum_{i=1}^{n}Z_{i}\to p_{z}$). Then, for $z\in\{0,1\}$ and $g\in\{1,\ldots,G\}$, if the following condition holds:
The results of Theorem A.2 can be viewed as a special case of Schochet (2024), Theorem 1. The condition in (6) is a Lindeberg-type condition, which restricts the tail behavior of the underlying potential outcome distribution and is relatively weak.
@@ A.5 Non-parametric test of treatment effect heterogeneity
[…]
where $\hat{\tau}:=\frac{1}{\sum_{i=1}^{n}Z_{i}}Y_{i}Z_{i}-\frac{1}{\sum_{i=1}^{n}(1-Z_{i})}Y_{i}(1-Z_{i})$ is the usual difference-in-means estimator across the entire sample. Then, following Theorem A.2, we can show that under the null hypothesis, the test statistic can be asymptotically approximated using a $\chi^{2}$ distribution. The following corollary formalizes.
@@
$\boldsymbol{\hat{\tau}}^{\top}\boldsymbol{\Sigma}^{-1}\boldsymbol{\hat{\tau}}\stackrel{{\scriptstyle d}}{{\rightarrow}}\chi^{2}.$
@@ A.6 Detailed CDT Algorithm
2 // Stage 1: estimate heterogeneous treatment effects via teacher model
@@ A.7 Discussion of Teacher Model Selection Algorithm
2   Fit $\mathcal{M}^{l}$ on $\mathcal{D}^{train}$ to obtain estimates of the heterogeneous treatment effects $\hat{\tau}^{d,l}_{i}$ for each unit in $\mathcal{D}^{train}$ (via lines 2-16 in Algorithm 1)
@@ Appendix B Proofs and Derivations
$\qquad\times\Pr\left[\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(a)},\hat{s}_{n}^{(a)})\right\}<\mathbb{E}_{n}\left\{\ell(\hat{\tau}_{i}^{d};X^{(b)},\hat{s}_{n}^{(b)})\right\}\right]$
@@
B.3 Proof of Theorem A.1
Proof.
To prove Theorem A.1, we will introduce a lemma that will be helpful in the following derivation. We define $n_{g}$ as the total number of units in a subgroup $g$, and $n_{g}(1)=\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})Z_{i}$ – i.e., $n_{g}(1)$ is equal to the number of units in subgroup $g$ who are treated. We similarly define $n_{g}(0)$ as the number of units in subgroup $g$ who are not treated. Throughout, we will denote $\mathbb{E}_{\mathcal{D}}(A_{i})=\mathbb{E}(A_{i}\mid\mathcal{D})$ as the expectation, given the sample data $\mathcal{D}$.
@@
    Equiprobable treatment assignment patterns. For each subgroup $g\in\{1,\ldots,G\}$, all $n_{g}\choose n_{g}(1)$ combination of ways to treat $n_{g}(1)$ units are equiprobable, given the subgroup size $n_{g}$.
@@
We are considering designs that satisfy assignment symmetry. For example, experimental designs such as complete randomization, Bernoulli randomization, or blocking with independent treatment assignment across blocks, would satisfy assignment symmetry. However, cluster randomization would not satisfy assignment symmetry. Furthermore, we do not account for settings in which some units within a subgroup are more likely to receive treatment.
@@
$\mathbb{E}_{\mathcal{D}}\left\{Z_{i}\mid n_{g}(1)\right\}=\frac{n_{g}(1)}{n_{g}}\equiv\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})Z_{i},\text{ and }\mathbb{E}_{\mathcal{D}}\left\{Z_{i}Z_{j}\mid n_{g}(1)=a\right\}=\frac{a(a-1)}{n_{g}(n_{g}-1)}$ (9)
@@
  • $\mathbb{E}_{\mathcal{D}}\left[\frac{Z_{i}}{n_{g}(1)}\right]=\frac{1}{n_{g}}$
  • $\mathbb{E}_{\mathcal{D}}\left[\frac{Z_{i}^{2}}{n_{g}(1)^{2}}\right]=\frac{1}{n_{g}}\mathbb{E}_{\mathcal{D}}\left[\frac{1}{n_{g}(1)}\right]$
  • $\mathbb{E}_{\mathcal{D}}\left[\frac{Z_{i}Z_{j}}{n_{g}(1)^{2}}\right]=\frac{1}{n_{g}(n_{g}-1)}\mathbb{E}_{\mathcal{D}}\left[\frac{n_{g}(1)-1}{n_{g}(1)}\right]=\frac{1}{n_{g}(n_{g}-1)}\left(1-\mathbb{E}_{\mathcal{D}}\left[\frac{1}{n_{g}(1)}\right]\right)$
  • $\mathbb{E}_{\mathcal{D}}\left[\frac{Z_{i}(1-Z_{j})}{n_{g}(1)n_{g}(0)}\right]=\frac{1}{n_{g}\cdot(n_{g}-1)}\cdot\mathbb{E}_{\mathcal{D}}\left[\frac{n_{g}-n_{g}(1)}{n_{g}(0)}\right]=\frac{1}{n_{g}\cdot(n_{g}-1)}$
@@
Part 1. Unbiasedness. To begin, we will show that $\hat{\tau}^{(g)}$ is unbiased for the within-sample subgroup ATE:
$\mathbb{E}_{\mathcal{D}}$ $\left(\hat{\tau}^{(g)}\right)=\mathbb{E}\left\{\left.\frac{1}{\sum_{i=1}^{n}Z_{i}\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}Z_{i}Y_{i}\hat{\mathcal{G}}_{g}(X_{i})-\frac{1}{\sum_{i=1}^{n}(1-Z_{i})\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}(1-Z_{i})Y_{i}\hat{\mathcal{G}}_{g}(X_{i})~\right|\mathcal{D}\right\}$
@@
$\text{var}_{\mathcal{D}}$ $\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1)\right\}$ (12)
$=\sum_{i,j:\hat{\mathcal{G}}_{g}(X)=1}\text{cov}_{\mathcal{D}}\left(\frac{Z_{i}}{n_{g}(1)}Y_{i}(1),\frac{Z_{j}}{n_{g}(1)}Y_{j}(1)\right)$
$=\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[\text{var}_{\mathcal{D}}\left\{\frac{Z_{i}}{n_{g}(1)}Y_{i}(1)\right\}+\sum_{i\neq j}\text{cov}_{\mathcal{D}}\left\{\frac{Z_{i}}{n_{g}(1)}Y_{i}(1),\frac{Z_{j}}{n_{g}(1)}Y_{j}(1)\right\}\right]$
$=\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[\frac{1}{n_{g}}\mathbb{E}_{\mathcal{D}}\left\{\frac{1}{n_{g}(1)}\right\}Y_{i}(1)^{2}-\frac{1}{n_{g}^{2}}Y_{i}(1)^{2}\right]+$
$\sum_{i\neq j}\frac{1}{n_{g}}\cdot\frac{1}{n_{g}-1}\cdot\left(1-\mathbb{E}_{\mathcal{D}}\left\{\frac{1}{n_{g}(1)}\right\}\right)Y_{i}(1)Y_{j}(1)-\frac{1}{n_{g}^{2}}Y_{i}(1)Y_{j}(1)$
$=\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left(\mathbb{E}_{\mathcal{D}}\left\{\frac{1}{n_{g}(1)}\right\}-\frac{1}{n_{g}}\right)Y_{i}(1)^{2}+$
$~~~~\frac{1}{n_{g}}\sum_{i\neq j:\hat{\mathcal{G}}_{g}(X)=1}\left(\frac{1}{n_{g}-1}\left(1-\mathbb{E}_{\mathcal{D}}\left\{\frac{1}{n_{g}(1)}\right\}\right)-\frac{1}{n_{g}}\right)Y_{i}(1)Y_{j}(1)$
Let $\tilde{\beta}_{g}(z):=\mathbb{E}_{\mathcal{D}}\{1/n_{g}(1)\}$:
@@
$=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\cdot\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[Y_{i}(1)^{2}-\sum_{i\neq j}Y_{i}(1)Y_{j}(1)\right]$
$=\left\{\tilde{\beta}_{g}(1)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(1)\mid\mathcal{D}\right\},$ (13)
where $\text{var}_{g}\left\{Y_{i}(1)\mid\mathcal{D}\right\}:=\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left[Y_{i}(1)^{2}-\sum_{i\neq j}Y_{i}(1)Y_{j}(1)\right]$. We can similarly show term (2) can be rewritten as:
$\text{var}_{\mathcal{D}}\left\{\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0)\right\}=\left\{\beta_{g}(0)-\frac{1}{n_{g}}\right\}\text{var}_{g}\left\{Y_{i}(0)\mid\mathcal{D}\right\}.$ (14)
@@
$\mathbb{E}_{\mathcal{D}}$ $\left[\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1)\right\}\left\{\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0)\right\}\right]$
$=\mathbb{E}_{\mathcal{D}}\left\{\frac{1}{n_{g}(1)\cdot n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}(1-Z_{i})Y_{i}(1)Y_{i}(0)+\sum_{i\neq j}Z_{i}(1-Z_{j})Y_{i}(1)Y_{j}(0)\right\}$
@@
$=\mathbb{E}_{\mathcal{D}}\left\{\frac{1}{n_{g}(1)\cdot n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\sum_{i\neq j}Z_{i}(1-Z_{j})Y_{i}(1)Y_{j}(0)\right\}$
@@
$\text{cov}_{\mathcal{D}}$ $\left\{\frac{1}{n_{g}(1)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}Z_{i}Y_{i}(1),\frac{1}{n_{g}(0)}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}(1-Z_{i})Y_{i}(0)\right\}$